OpenStudy (jkbo):

Please help Factor 25^2-64

OpenStudy (aonz):

you sure you got the question right?

OpenStudy (jkbo):

Yup thats the whole question

OpenStudy (aonz):

sure its not 25x^2 - 64?

HanAkoSolo (jamierox4ev3r):

\(25x^2-64\) @jkbo

OpenStudy (jkbo):

yes @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

oh ok so @AonZ was right then

OpenStudy (mayankdevnani):

\[\huge 25^2=625\] so we get, \[\huge 625-64\] and we can write this as : \[\huge (25)^2-(8)^2\] use \(\large a^2-b^2\) identity and this identity states that: \[\huge a^2-b^2=(a-b)(a+b)\]

OpenStudy (mayankdevnani):

can you solve it? @jkbo

OpenStudy (jkbo):

yes thank you I found the answer can you help me with one more? factor 15x^2 + 20x

OpenStudy (mayankdevnani):

@jkbo we can write this in two forms: \[\huge 1)~~~~ x(15x+20)\] \[\huge 2)~~~~5x(3x+4)\]

OpenStudy (mayankdevnani):

understood? @jkbo

OpenStudy (aonz):

well you would always want it it form 2 not 1

OpenStudy (mayankdevnani):

what? i don't understand. @AonZ

OpenStudy (aonz):

form 2 is more correct?

OpenStudy (mayankdevnani):

lol... how would you choose form 2 ? it is impossible . because in 20x , x is in the form of 1 degree not 2 degree How would we factorize it in 2 nd form?

OpenStudy (mayankdevnani):

@AonZ

OpenStudy (aonz):

Huh? your confusing me now?

OpenStudy (mayankdevnani):

lol..... can you factorize it in 2nd form \[\huge 20x\] can you?

OpenStudy (aonz):

oh, i meant when u said "we can write this in two forms" then I said number 2 is more correct

OpenStudy (mayankdevnani):

oh!!! all are correct

OpenStudy (jkbo):

15x +20

OpenStudy (mayankdevnani):

what is this? ^^^

OpenStudy (jkbo):

Thank you

OpenStudy (jkbo):

I was solving the problem

OpenStudy (mayankdevnani):

Ohkk

OpenStudy (jkbo):

Thank you all

OpenStudy (mayankdevnani):

welcome :)